OpenStudy (anonymous):

Which ordered pair does not satisfy the system 2x + y > 0 and -2x - y > -8? A. (3, -2) B. (2, 3) C. (-1, 4) D. (-1, -1)

OpenStudy (barrycarter):

For (A), x=3 and y= -2, so 2x+y=4 and -2x-y= -4. Since 4>0 and -4>-8 this one's OK. For (B), x=2 and y=3, so 2x+y=7 and -2x-y= -7. Again 3>0 and -7 > -8 so this one's one OK. For (C) x= -1, y=4, so 2x+y= 2 and -2x-y= - 2. Since 2>0 and -2>-8 this one's OK too. For (D) x=-1 and y=-1, so 2x+y=-3 ... That's NOT greater than zero, so this one doesn't fit.

OpenStudy (anonymous):

just plug each order pairs for x and y and see when it is not satisfy both inequalities